# CYCle: Choosing Your Collaborators Wisely to Enhance Collaborative Fairness in Decentralized Learning

### v1

@@ 5.2 Experimental Results
We start with benchmarking existing algorithms on the CIFAR-10 dataset. Table 1 summarizes the results for different splitting strategies involving $N=5$ participants. Our key findings are: (i) We assess the performance of three algorithms - FedAvg, VPDL, and CYCle - using the ResNet18 architecture. These results are shown in Figure 2 and the upper section of Table 1 (rows 3-5). FedAvg has better MCG compared to other methods due to its centralized FL approach, which involves aggregating and sharing full model parameters, resulting in good overall utility. However, when it comes to collaborative fairness, FedAvg falls short of being fair to participants who contribute a larger amount of data. In particular, the CYCle approach exhibits significantly lower CGS values compared to the FedAvg algorithm in the imbalanced settings. Additionally, our method consistently provides positive collaboration gain to all participants across varied settings. This is not the case for FedAvg or VPDL, which is evident in the outcomes for $\mathcal{P}_{1}$ in the imbalanced $(0.8,1)$ and $(0.6,1)$ settings. (ii) For comparison with other fairness-aware FL algorithms, we utilize a custom CNN architecture used in [28]. This is necessary because CGSV works better on simpler architectures. The results are shown in Figure 3 and the bottom part of Table 1 (rows 6-9). Our method has better or comparable CG values when compared to CFFL, RFFL, and CGSV in most of the scenarios. In terms of fairness, we surpass other approaches by achieving the lowest CGS values in all but one scenario: imbalanced $(0.8,1)$ setting, where it is on par with RFFL. Notably, CYCle consistently achieves positive gains for all participants, whereas other methods degrade performance for the most contributing participant (see Fig. 3).
Table 2: Collaboration gain and spread (MCG, CGS) on CIFAR-10 and CIFAR-100 datasets and the given partition strategies for $N=\{5,10\}$ using our proposed algorithm.
@@
Next, we assess the performance of our approach on the CIFAR-10 and CIFAR-100 datasets with different number of participants ($N\in\{5,10\}$) and data splitting scenarios (Tab. 2). In all the settings, we consistently achieve positive collaboration gain for all participants and the CYCle method consistently has lower CGS compared to VPDL, indicating the effectiveness of our method.
@@
Reputation scoring visualization: Fig. 4 visualizes the computed reputation scores in three collaboration rounds: $t=0,25,75$ using heatmap plots. In Fig 4(a), where participants have imbalanced data (with $\mathcal{P}_{1}$ having 80% of the data), our method can already identify distinct patterns of each participant’s data value from the beginning of the collaboration. The heatmap figure emphasizes that the higher is importance given to $\mathcal{P}_{1}$ by all the other participants. As the collaboration progresses, the reputation of $\mathcal{P}_{1}$ intensifies further, while the reputation scores of other participants decrease. Towards the end of the collaboration, the values converge to a state where participants with an equal amount of data receive equal reputation scores. A similar pattern can be observed in Fig 4(b), where the first two participants possess larger amounts of data. While the first five rows of the heatmap figures depict the raw reputation scores calculated using Eq. 5, the sixth row at the bottom represents the average reputation scores aggregated over all the rounds.
Figure 5: Heatmap visualization of reputation scores when the labels at one of the participants $\mathcal{P}_{5}$ get corrupted by label flipping. When the flip rate is large, the “malicious” client is quickly identified by the other participants and it gets shunted out of the collaboration.
@@
Free Rider Study: In this study involving $N=5$ participants with homogeneous data splitting scenario, we simulate varying degrees of label flipping at participant $\mathcal{P}_{5}$. We tested the label flipping rates of $0.2$, $0.5$ and $1.0$. From Fig. 5, it is evident that the honest participants (participants $\mathcal{P}_{1},...,\mathcal{P}_{4}$) quickly discern the anomaly with $\mathcal{P}_{5}$, detecting it as a potential threat. Consequently, they assign a reputation score of 0.0 to $\mathcal{P}_{5}$, indicating a cessation of updates to this participant. On the other hand, $\mathcal{P}_{5}$ accurately recognized that the other participants were beneficial when the label flipping rate was at $0.2$, but this positive assessment drops to zero for the case when the label space of $\mathcal{P}_{5}$ become completely different.
@@ 6 Conclusion
This paper proposes a novel reputation-based adaptive sharing algorithm, specifically tailored for decentralized learning scenarios. Through theoretical and empirical analyses, this work has illustrated the algorithm’s superior capabilities in achieving collaborative fairness. In future, we will theoretically study the convergence characteristics of the proposed protocol.
@@ Appendix A Mean Estimation under CYCle Protocol
Standard FL. The vanilla FedAvg algorithm, a standard FL approach, does not ensure an optimal final model for all participants in a federated setting. Consider a simple scenario involving two clients, $N=2$, each aiming to estimate the mean of their data distribution by minimizing the loss function $f_{k}(w)=(w-\theta_{k})^{2}$. However, these clients can’t compute the true mean directly due to having only $N_{k}$ samples from their distributions, denoted $e_{k,j}\sim\mathcal{N}(\theta_{k},\nu^{2}),\forall j\in[N_{k}]$. Consequently, they minimize the empirical loss function $F_{k}(w)=(w-\widehat{\theta}_{k})^{2}+(\widehat{\theta}_{k}-\theta_{k})^{2}$, where $\widehat{\theta}_{k}=\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}e_{k,j}$, leading to $\widehat{w}_{k}=\widehat{\theta}_{k}$ as the minimizer of $F_{k}(w)$.
@@
Note that the distribution of the empirical means themselves are distributed normally, due to the linear additivity property of independent normal random variables. This setting is similar to [6, 5], albeit with a few modifications.
@@
(Cho et al. [6]) The probability (likelihood) that the model obtained from collaborative training outperforms the standalone training model is upper-bounded by $2\exp\Big{(}-\dfrac{\gamma_{G}^{2}}{5\gamma^{2}}\Big{)}$ in a standard FL framework.
@@
=($=\mathbb{P}\Bigg{(}\Bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}\leq 0\Bigg{\}}\cap\Big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\Big{\}}\Bigg{)}$
+($\qquad+\mathbb{P}\Bigg{(}\Bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}\leq 0\Bigg{\}}\cap\Big{\{}\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\Big{\}}\Bigg{)}$ (12)
=($=\mathbb{P}\Bigg{(}\Bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\Bigg{\}}\cap\Big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\Big{\}}\Bigg{)}$
+($\qquad+\mathbb{P}\Bigg{(}\Bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}\leq 0\Bigg{\}}\cap\Big{\{}\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\Big{\}}\Bigg{)}$ (13)
@@
CYCle Algorithm. The CYCle algorithm uniquely incorporates the reputations of each client during the aggregation of their computed means. Considering the scenario from the perspective of Client 1, who possesses an estimate $\widehat{\theta}_{1}$ and receives $\widehat{\theta}_{2}$ from Client 2, the aggregation is expressed as:
@@
Reputation Scoring. In the CYCle algorithm, each participant calculates the reputation of their collaborators by assessing the distance between their shared means, denoted as $d=\Big{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\Big{)}^{2}$. This distance is then used to determine their reputations as follows:
@@
Cost Function. To evaluate the CYCle algorithm against the standard, centralized FL algorithm, we consider the empirical losses each participant ends up solving. In FedAvg, all clients achieve the common mean:
@@
Minimizing the CYCle Objective: Ensuring Positive Outcomes. Unlike standard FL where increasing true heterogeneity $\gamma_{G}^{2}$ drastically reduces the usefulness of the global model (as noted in Lemma A.1), CYCle modulates the aggregation phase using reputation scores which are derived from the empirical heterogeneity parameter $\widehat{\gamma}_{G}^{2}$ (19).
Theorem A.1.
The probability (likelihood) that the model obtained through collaboration ($w$) surpasses a standalone-trained model $(\widehat{\theta})$ is lower bounded by $\dfrac{1}{8}\exp{\big{(}-\dfrac{1}{4\gamma^{2}}\big{)}}$ in the CYCle algorithm.
@@
=($=\mathbb{P}\Bigg{(}\bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}^{2}+2\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}(\widehat{\theta}_{1}-\theta_{1})\leq 0\bigg{\}}\cap\big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\big{\}}\Bigg{)}$
+($\qquad+\mathbb{P}\Bigg{(}\bigg{\{}\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}^{2}+2\bigg{(}\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg{)}(\widehat{\theta}_{1}-\theta_{1})\leq 0\bigg{\}}\cap\big{\{}\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\big{\}}\Bigg{)}$ (28)
($\geq\mathbb{P}\Big{(}\big{\{}1+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\big{\}}\cap\big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\big{\}}\Big{)}$ (29)
=($=\mathbb{P}\Big{(}\big{\{}\widehat{\theta}_{1}-\theta_{1}\leq-\frac{1}{2}\big{\}}\cap\big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\big{\}}\Big{)}$ (30)
@@
where (27) uses the reputation score $r=1$ when $\widehat{\gamma}_{G}^{2}<1$ (refer to (19)), (28) uses $\mathbb{P}(A)=\mathbb{P}(A\cap B)+\mathbb{P}(A\cap B^{C})$, (29) uses the definition of $\widehat{\gamma}_{G}^{2}\leq 1$ and the second part of expression can be dropped since $\widehat{\gamma}_{G}$ can take values from $-1$ to 00, it can be assigned zero probability, (32) uses $\mathbb{P}(A|B)\geq\mathbb{P}(A)$, (33) uses $\widehat{\theta}_{1}-\theta_{1}\sim\mathcal{N}(0,\gamma^{2})$, (34) $\mathbb{P}(\widehat{\theta}_{1}<\widehat{\theta}_{2})\geq\frac{1}{2}$ and $\mathbb{P}(Z\geq x)\geq\frac{1}{4}\exp(-x^{2})$ where $Z\sim\mathcal{N}(0,1)$. The similar bound can be found for the second client.
@@
($\geq\mathbb{P}\Big{(}\big{\{}1+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\big{\}}\cap\big{\{}\widehat{\theta}_{2}>\widehat{\theta}_{1}\big{\}}\Big{)}$ (39)
@@
where $w_{1}=\Big{(}1-\dfrac{r}{2}\Big{)}\widehat{\theta}_{1}+\dfrac{r}{2}\widehat{\theta}_{2}=\widehat{\theta}_{1}$, since $r=0$ for $\widehat{\gamma}_{G}^{2}>2$ (Eqs. 22, 19).

### v2

@@ 5.2 Experimental Results
We start by benchmarking existing algorithms on the CIFAR-10 dataset. Table 5.1 summarizes the results for different splitting strategies that involve $N=5$ participants. Our key findings are as follows:
  • We assess the performance of three algorithms – FedAvg, VPDL, and CYCle – using the ResNet18 architecture. These results are shown in Figure 3 and in the upper section of Table 5.1 (rows 3-5). FedAvg has better MCG compared to other methods due to its centralized FL approach, which involves aggregating and sharing full model parameters, resulting in good overall utility. However, when it comes to collaborative fairness, FedAvg falls short of being fair to participants who contribute a larger amount of data. In particular, the CYCle approach exhibits significantly lower CGS values compared to the FedAvg algorithm in the imbalanced settings. Additionally, our method consistently provides positive collaboration gains to all participants in various settings. This is not the case for FedAvg or VPDL, which is evident in the outcomes for $\mathcal{P}_{1}$ in the imbalanced $(0.8,1)$ and $(0.6,1)$ settings.
  • For comparison with other fairness-aware FL algorithms, we utilize a custom CNN architecture used in Xu et al. (2021). This is necessary because CGSV works better on simpler architectures. The results are shown in Fig. 4 and the bottom part of Tab. 5.1 (rows 6-9). Our method has better or comparable CG values when compared to CFFL, RFFL, and CGSV in most of the scenarios. In terms of fairness, we surpass other approaches by achieving the lowest CGS values in all but one scenario: the imbalanced $(0.8,1)$ setting, where it is on par with RFFL. Notably, CYCle consistently achieves positive gains for all participants, whereas other methods degrade performance for the most contributing participant (see Figure 4).
Figure 4: Per-participant performance comparison on CIFAR-10 dataset using custom CNN architecture from (Xu et al., 2021): Validation accuracy evaluated with $N=5$ participants.
Table 2: Collaboration gain and spread (MCG, CGS) on CIFAR-10 and CIFAR-100 datasets and the given partition strategies for $N=\{5,10\}$ using our proposed algorithm. DSGD refers to the Gossip-SGD algorithm.
@@
DSGD $83.72$ $10.02$ $0.56$ $80.38$ $15.65$ $1.17$ $56.06$ $20.02$ $0.54$ $50.33$ $24.94$ $0.67$
CYCle $83.85$ $10.15$ $0.33$ $80.46$ $15.73$ $0.77$ $55.87$ $19.83$ $0.48$ $49.62$ $24.24$ $0.43$
Next, we assess the performance of our approach on the CIFAR-10 and CIFAR-100 datasets with different numbers of participants ($N\in\{5,10\}$) and data splitting scenarios (Table 2). In all settings, we consistently achieve a positive collaboration gain for all participants, and the CYCle method consistently has lower CGS compared to VPDL, demonstrating improved fairness across participants. Furthermore, we observe that when integrated with Gossip-SGD (DSGD), CYCle maintains this trend: while the baseline DSGD achieves high MCG values due to aggressive mixing, it often results in large CGS values, indicating significant disparity among participants. In contrast, CYCle significantly reduces CGS without compromising MCG, highlighting its ability to regulate knowledge flow more equitably in decentralized and peer-to-peer settings.
(a) Imbalanced (0.8, 1) setting
(b) Imbalanced (0.35, 2) setting
Reputation scoring visualization.
@@
Figure 5 visualizes the reputation scores computed in three collaboration rounds: $t=0,25,75$ using heatmap plots. In Figure 5(a), where participants have imbalanced data (with $\mathcal{P}_{1}$ having 80% of the data), our method can already identify distinct patterns of each participant’s data value from the beginning of the collaboration. The heatmap figure emphasizes that higher importance is given to $\mathcal{P}_{1}$ by all the other participants. As the collaboration progresses, the reputation of $\mathcal{P}_{1}$ intensifies further, while the reputation scores of other participants decrease. Toward the end of the collaboration, the values converge to a state in which participants with an equal amount of data receive equal reputation scores. A similar pattern can be observed in Figure 5(b), where the first two participants possess larger amounts of data. While the first five rows of the heatmap figures depict the raw reputation scores calculated using Equation 7, the sixth row at the bottom represents the average reputation scores aggregated over all the rounds.
Standalone ($\mathcal{B}_{n}$) $83.42$ $83.67$ $83.15$ $83.13$ $83.63$ $83.34$
Flip rate = 0.0 $86.12$ $87.03$ $86.33$ $85.62$ $86.10$ $86.28$
Flip rate = 0.2 $84.95$ $84.5$ $84.5$ $86.45$ $60.95$ $85.10$
Flip rate = 0.5 $84.25$ $84.65$ $84.65$ $83.25$ $55.80$ $84.20$
Flip rate = 1.0 $83.9$ $84.5$ $83.95$ $85.1$ $10.75$ $84.36$
Figure 6: Heatmap of reputation scores when $\mathcal{P}_{5}$ undergoes label flipping. Higher flip rates lead to $\mathcal{P}_{5}$ being rapidly identified as malicious by others.
Figure 7: Per-participant accuracy under CYCle (Fed-ISIC2019).
Free Rider Study.
@@
In this study involving $N=5$ participants with a homogeneous data splitting scenario, we simulate varying degrees of label flipping at participant $\mathcal{P}_{5}$. We tested the label flipping rates of $0.2$, $0.5$, and $1.0$. From Figure 6, it is evident that honest participants (participants $\mathcal{P}_{1},...,\mathcal{P}_{4}$) quickly discern the anomaly with $\mathcal{P}_{5}$, detecting it as a potential threat. Consequently, they assign a reputation score of 0.0 to $\mathcal{P}_{5}$, indicating a cessation of updates to this participant. On the other hand, $\mathcal{P}_{5}$ accurately recognized that the other participants were beneficial when the label flipping rate was at $0.2$, but this positive assessment drops to zero for the case when the label space of $\mathcal{P}_{5}$ becomes completely different.
Table 6, aligned with Figure 6, details the values of $\mathcal{B}_{n}$ and $\mathcal{A}_{n}$. It reveals that, in the presence of flipped labels, the collaboration gain declines, reflecting a loss of a participant that holds 20% of the data.
Table 4: CYCle vs. Gossip-SGD on MCG/CGS (Fed-ISIC2019).
Method MCG CGS
Gossip-SGD (Complete) 30.26 15.64
CYCle 23.79 10.94
Evaluation on Real-World Medical Dataset (Fed-ISIC2019).
To validate our framework in real-world heterogeneous settings, we evaluate CYCle on the non-IID Fed-ISIC2019 dataset (Ogier du Terrail et al., 2022). This benchmark features complex inter-client shifts in class distributions and image types.
Figure 7 shows the per-client validation accuracy using our CYCle extension on top of Gossip-SGD. Despite differences in client data sizes and class coverage, CYCle maintains high fairness across participants, as evidenced by a strong correlation between client contribution and reward ($\rho=0.9577$). Table 4 confirms that CYCle achieves the lowest CGS while preserving collaborative utility.
The complete experimental setup, alternate topologies, and detailed visualizations (including reputation dynamics and data distribution) are provided in Appendix G.3.
@@ 6 Conclusion
This paper proposed a novel reputation-based adaptive sharing algorithm to promote collaborative fairness, specifically tailored for decentralized learning scenarios. By leveraging gradient alignment, CYCle dynamically adjusts knowledge transfer based on participants’ contributions, ensuring positive and equitable collaboration gains. We further extended the protocol to operate seamlessly over gossip-based decentralized learning algorithms. Through theoretical and empirical analyses, this work has illustrated the algorithm’s superior capabilities in achieving collaborative fairness.
Broader Impact Statement
Privacy Protection Measures.
CYCle inherits the secure knowledge-sharing protocol of the CaPriDe learning framework (Tastan & Nandakumar, 2023), which uses fully homomorphic encryption (FHE) to safeguard sensitive data. Specifically, each client only transmits encrypted prediction logits to peers, never raw data (e.g. medical images from Fed-ISIC2019) or model parameters. These encrypted outputs are used solely to compute a local distillation loss on peer-held datasets, and they are never decrypted by any party. This ensures that all sensitive information remains protected throughout the training process.
Abuse Risk and Mitigation.
A common concern in collaborative learning is the potential for knowledge monopolization, where dominant participants attempt to suppress or filter contributions from less performant peers. CYCle directly addresses this risk via its similarity-based, bidirectional collaboration mechanism. Specifically, clients benefit most when they are aligned with others (as defined by gradient similarity), which discourages unilateral exclusion. Furthermore, the soft contribution mapping function $h(s)$ makes it suboptimal for any client to reject collaboration entirely, as doing so reduces their own update quality. This design promotes a balanced and inclusive exchange of knowledge.
Acknowledgments
This material is partly based on work supported by the Office of Naval Research N00014-24-1-2168.
@@ Appendix A Mean Estimation under CYCle Protocol
Standard FL.
The vanilla FedAvg algorithm, a standard FL approach, does not ensure an optimal final model for all participants in a federated setting. Consider a simple scenario involving two clients, $N=2$, each aiming to estimate the mean of their data distribution by minimizing the loss function $f_{k}(w)=(w-\theta_{k})^{2}$. However, these clients can’t compute the true mean directly due to having only $N_{k}$ samples from their distributions, denoted $e_{k,j}\sim\mathcal{N}(\theta_{k},\nu^{2}),\forall j\in[N_{k}]$. Consequently, they minimize the empirical loss function $F_{k}(w)=(w-\widehat{\theta}_{k})^{2}+(\widehat{\theta}_{k}-\theta_{k})^{2}$, where $\widehat{\theta}_{k}=\frac{1}{N_{k}}\sum_{j=1}^{N_{k}}e_{k,j}$, leading to $\widehat{w}_{k}=\widehat{\theta}_{k}$ as the minimizer of $F_{k}(w)$.
@@
Note that the distribution of the empirical means themselves is distributed normally, due to the linear additivity property of independent normal random variables. This setting is similar to (Cho et al., 2022; 2024), albeit with a few modifications.
@@
(Cho et al., 2022) The probability (likelihood) that the model obtained from collaborative training outperforms the standalone training model is upper-bounded by $2\exp\Big(-\dfrac{\gamma_{G}^{2}}{5\gamma^{2}}\Big)$ in a standard FL framework.
@@
$=\mathbb{P}\Bigg(\Bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)\leq 0\Bigg\}\cap\Big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\Big\}\Bigg)$
$\qquad+\mathbb{P}\Bigg(\Bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)\leq 0\Bigg\}\cap\Big\{\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\Big\}\Bigg)$ (15)
$=\mathbb{P}\Bigg(\Bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\Bigg\}\cap\Big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\Big\}\Bigg)$
$\qquad+\mathbb{P}\Bigg(\Bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)^{2}+2(\widehat{\theta}_{1}-\theta_{1})\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)\leq 0\Bigg\}\cap\Big\{\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\Big\}\Bigg)$ (16)
@@
CYCle Algorithm.
The CYCle algorithm uniquely incorporates the reputations of each client during the aggregation of their computed means. Considering the scenario from the perspective of Client 1, who possesses an estimate $\widehat{\theta}_{1}$ and receives $\widehat{\theta}_{2}$ from Client 2, the aggregation is expressed as:
@@
Reputation Scoring.
In the CYCle algorithm, each participant calculates the reputation of their collaborators by assessing the distance between their shared means, denoted as $d=\Big(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\Big)^{2}$. This distance is then used to determine their reputations as follows:
@@
Cost Function.
To evaluate the CYCle algorithm against the standard, centralized FL algorithm, we consider the empirical losses each participant ends up solving. In FedAvg, all clients achieve the common mean:
@@
Minimizing the CYCle Objective: Ensuring Positive Outcomes.
Unlike standard FL where increasing true heterogeneity $\gamma_{G}^{2}$ drastically reduces the usefulness of the global model (as noted in Lemma A.1), CYCle modulates the aggregation phase using reputation scores which are derived from the empirical heterogeneity parameter $\widehat{\gamma}_{G}^{2}$ (22).
Theorem A.1 (Restated Theorem 4.1).
[…]
@@
$=\mathbb{P}\Bigg(\bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)^{2}+2\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)(\widehat{\theta}_{1}-\theta_{1})\leq 0\bigg\}\cap\big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\big\}\Bigg)$
$\qquad+\mathbb{P}\Bigg(\bigg\{\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)^{2}+2\bigg(\frac{\widehat{\theta}_{2}-\widehat{\theta}_{1}}{2}\bigg)(\widehat{\theta}_{1}-\theta_{1})\leq 0\bigg\}\cap\big\{\widehat{\theta}_{2}\leq\widehat{\theta}_{1}\big\}\Bigg)$ (31)
$\geq\mathbb{P}\Big(\big\{1+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\big\}\cap\big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\big\}\Big)$ (32)
$=\mathbb{P}\Big(\big\{\widehat{\theta}_{1}-\theta_{1}\leq-\frac{1}{2}\big\}\cap\big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\big\}\Big)$ (33)
@@
where (30) uses the reputation score $r=1$ when $\widehat{\gamma}_{G}^{2}<1$ (refer to (22)), (31) uses $\mathbb{P}(A)=\mathbb{P}(A\cap B)+\mathbb{P}(A\cap B^{C})$, (32) uses the definition of $\widehat{\gamma}_{G}^{2}\leq 1$ and the second part of the expression can be dropped since $\widehat{\gamma}_{G}$ can take values from 1-1 to 0, it can be assigned zero probability, (35) uses $\mathbb{P}(A|B)\geq\mathbb{P}(A)$, (36) uses $\widehat{\theta}_{1}-\theta_{1}\sim\mathcal{N}(0,\gamma^{2})$, (37) $\mathbb{P}(\widehat{\theta}_{1}<\widehat{\theta}_{2})\geq\frac{1}{2}$ and $\mathbb{P}(Z\geq x)\geq\frac{1}{4}\exp(-x^{2})$ where $Z\sim\mathcal{N}(0,1)$. The similar bound can be found for the second client.
@@
$\geq\mathbb{P}\Big(\big\{1+2(\widehat{\theta}_{1}-\theta_{1})\leq 0\big\}\cap\big\{\widehat{\theta}_{2}>\widehat{\theta}_{1}\big\}\Big)$ (42)
@@
where $w_{1}=\Big(1-\dfrac{r}{2}\Big)\widehat{\theta}_{1}+\dfrac{r}{2}\widehat{\theta}_{2}=\widehat{\theta}_{1}$, since $r=0$ for $\widehat{\gamma}_{G}^{2}>2$ (Eqs. 25, 22).
@@
Scope and Limitations.
The result in Theorem 4.1 is derived under a simplified two-client mean estimation setting. While this toy model is useful to illustrate how collaboration may fail under heterogeneity and how CYCle can mitigate that through similarity-based reputation, it is not intended as a universal lower bound for all learning scenarios. In particular, the lower bound on success probability does not generalize trivially to more complex settings (e.g., neural networks, high-dimensional tasks, or $N>2$ clients).
Extending the analysis to larger populations introduces new technical challenges. Specifically, when $N>2$, (i) the reputation becomes a full matrix rather than a scalar, (ii) true-mean gaps form a vector rather than a single $\gamma_{G}$, and (iii) the same noisy estimate $\widehat{\theta}_{i}$ appears in multiple clients’ aggregates, introducing cross-correlations. The bias term then depends on the worst-case alignment across all pairs, while the variance term includes complex cross-terms from covariance expansion. Bounding these quantities is non-trivial and remains an open problem.
We include this example primarily to explain the mechanism by which data heterogeneity undermines fairness in standard FL approaches, and to motivate our design choices in CYCle. Developing more general guarantees is an exciting direction for future work.